Solve by Substitution method
9m+n=32
m-3n=16

Answer by mananth(16949)

9m+n=32...........1
m-3n=16...........2
..
multiply equation by 3 and it to equation 2
3(9m+n)+(m-3n)=32*3+16
27m+3n+m-3n=96+16
28m=112
divide by 28
28m/28=112/28
m=4
..
plug the value of m in the equation 2
4-3n=16
Add -4 on both sides
4-4-3n=16-4
-3n=12
divide by -3
-3n/-3 = 12/-3
n=-4